1. A researcher estimates that the average height of the buildings of 30 or more stories in a large city is at least 700 feet. A random sample of 10 buildings is selected, and the heights in feet are shown. At  = 0.025, is there enough evidence to reject the claim? 485 511 841 725 615 520 535 635 616 582


1
Expert's answer
2022-06-08T16:15:05-0400
"\bar{x}=\dfrac{1}{10}(485 +511+ 841+ 725+ 615"

"+520 +535 +635 +616+ 582)=606.5"

"s^2=\dfrac{\sum_i(x_i-\bar{x})^2}{n-1}=\dfrac{1}{10-1}((485-606.5)^2"

"+(511-606.5)^2+(841-606.5)^2"

"+(725-606.5)^2+(615-606.5)^2"

"+(520-606.5)^2+(535-606.5)^2"

"+(635-606.5)^2+(616-606.5)^2"

"+(582-606.5)^2)=\dfrac{107084.5}{9}"

"s=\sqrt{s^2}=\sqrt{\dfrac{107084.5}{9}}\approx109.08"

The parameter is average content of fruit concentrate per bottle.

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge700"

"H_1:\mu<700"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.025," "df=n-1=9" and the critical value for a left-tailed test is "t_c =-2.262156."

The rejection region for this left-tailed test is "R = \{t:t<-2.262156\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{606.5-700}{109.08/\sqrt{10}}=-2.7106"


Since it is observed that "t=-2.7106<-2.262156=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for left-tailed, "df=9" degrees of freedom, "t=-2.7106" is "p=0.011987," and since "p=0.011987<0.025=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is less than 700, at the "\alpha = 0.025" significance level.